# Incentive-Aligned Multi-Source LLM Summaries

### v1

@@ 4.2 Result 1: Robustness against adversarial and untruthful sources
As a result of this clear separation between reliable and unereliable sources, we are able to see significant improvement gain in accuracy for both the NQ and ClashEval dataset in Table 1 and 2. Fluency also improves: see App. I.1 (Table 3).
@@
Initial Synthesis 38.3% 20.7% 26.9% 68/300 (22.7%)
Majority Prompt 39.6% 20.0% 26.6% 73/300 (24.3%)
Majority Claims 44.6% 19.8% 27.4% 102/300 (34.0%)
Our Method (TTS) 81.0% 31.9% 45.7% 212/300 (70.7%)
@@
Initial Synthesis 39.6% 16.8% 23.6% 10/300 (3.3%)
Majority Prompt 48.6% 21.3% 29.7% 19/300 (6.3%)
Majority Claims 46.3% 16.0% 23.8% 42/300 (14.0%)
Our Method (TTS) 86.4% 26.4% 40.4% 223/300 (74.3%)
@@
Figure 3: Score separation and incentives. Left: Informative-agreement scores separate reliable from unreliable sources without labels. Right: Truthful behavior maximizes score against deviations in stance.
Incentive alignment in practice.
@@ 4.3 Result 2: Robustness against coordinated, uninformative behavior
One of the main advantages of the adapted multi-task peer prediction scoring rule is its robustness to coordinated, uninformative behavior, a canonical failure mode for simpler consensus-based systems. We test this in the ClashEval dataset by introducing a bloc of four “uninformative” sources strategically authored to contradict every claim. As shown in Figure 4, the naive majority-based scoring fails catastrophically. It not only rewards the colluding dummy sources, but as a byproduct, this pollution of the peer pool also falsely elevates the score of the adversarial source, causing it to be ranked higher than the genuinely truthful documents. In contrast, TTS correctly assigns near-zero scores to the uninformative bloc and robustly preserves the correct reliability ranking. More details are given in Appendix I.4.
@@
Takeaways.
Our experiments highlight two complementary advantages of the TTS framework. First, the two-pass pipeline provides structural robustness; by using a leave-one-out method to objectively score and filter unreliable sources before the final synthesis, it significantly improves the summary’s factual accuracy and fluency in a way that is robust to strategic attacks. Second, the informative-agreement score provides further incentive robustness. It rewards beyond-chance corroboration over raw consensus, allowing the mechanism to resist coordinated, uninformative strategies and correctly identify reliable sources without ground-truth labels. These empirical findings are consistent with our theoretical guarantees, demonstrating that the TTS framework makes truthful, careful reporting the most effective strategy for a source to be included in the final summary.
@@ Appendix A Related Works
Commercial search has already shifted toward LLM-written overviews that synthesize multiple pages (Google’s AI Overviews; Microsoft’s Copilot Search in Bing, Perplexity AI). In these experiences, citations are shown but the LLM determines salience and framing, moving competition from ranked links to representation in the overview itself (Google Search Blog, ; Bing Team, ; Perplexity AI, ).
@@
Our setting—multiple web sources of uneven quality, possibly in conflict—aligns with many research papers in the RAG domain. First, recent benchmarks systematize how to stress-test RAG beyond vanilla QA: they evaluate robustness to noise, counterfactuals, and long-context alternatives; provide explainable testbeds and failure analytics; and introduce standard tooling to compare systems (Chen et al., 2024; Friel et al., 2024; Rau et al., 2024; Li et al., 2025). Building on such evaluations, a second line studies how models arbitrate conflicts between internal priors and external evidence: ClashEval shows that state-of-the-art LLMs frequently adopt incorrect retrieved content over correct priors under controlled perturbations (Wu et al., 2024); subsequent methods reason explicitly over disagreement, e.g., AstuteRAG which elicits parametric knowledge, clusters internal/external evidence into consistent vs. conflicting sets, and finalizes answers by reliability (Wang et al., 2024), and MADAM-RAG which assigns each document to an agent, debates, and aggregates, evaluated on the RAMDocs dataset with ambiguity, misinformation, and noise (Wang et al., 2025). However, not all search scenarios should (or can) rely on internal priors of LLMs—for breaking news and evolving events, priors are stale. Consistent with our Introduction, we focus on settings where we either omit priors or treat them as just one more source when helpful. A third line introduces self-monitoring and corrective control: Self-RAG learns when to retrieve and to critique its own generations via reflection tokens (Asai et al., 2024), while Corrective RAG (CRAG) adds a retrieval-quality evaluator that triggers fallback actions (broaden web search, decompose/recompose) when evidence appears unreliable (Yan et al., 2024). Finally, when retrieval itself is corrupted, recent work documents how “blocker” or misleading documents can drive RAG below non-RAG baselines (Zeng et al., 2025) and develops defenses that isolate per-passage influence and certifiably d bound the impact of a limited number of corrupted contexts (Xiang et al., 2024). Complementary audits quantify how small amounts of synthetic misinformation materially degrade knowledge-intensive QA (Pan et al., 2023). These techniques harden fixed pipelines; by contrast, our goal is to reshape incentives so truthful reporting is the best strategy for sources.
@@
Because summaries now mediate attention, sources adapt to whatever the system rewards. Beyond classical prompt-injection via web content (Greshake et al., 2023), Nestaas et al. (2024) study so-called adversarial search engine optimization (SEO)—deliberately crafting pages to make an LLM favor them regardless of factual merit—including preference-manipulation attacks demonstrated against production LLM search and plugin ecosystems. Reports of hidden instructions in scholarly submissions targeting LLM-assisted review illustrate similar gaming incentives (Gibney, 2025). Our approach aims to dissuade such user-unfriendly manipulation by changing how sources are scored and fed into the summary.
@@
Incentive-aligned elicitation without verifiable truth is the province of peer prediction. Foundations include the Peer-Prediction method (Miller et al., 2005) and Bayesian Truth Serum (BTS) (Prelec, 2004), with robust BTS variants that work in small populations and for non-binary or continuous signals (Witkowski and Parkes, 2012; Radanovic and Faltings, 2013). Multi-task mechanisms address effort and uninformative agreement: output-agreement–style rules and their refinements establish strong or informed truthfulness given structure on signals (Dasgupta and Ghosh, 2013; Shnayder et al., 2016). Of particular relevance is Correlated Agreement (CA), which rewards informative (surprising) agreement across tasks rather than raw consensus; extensions handle heterogeneous tasks and heterogeneous user types, and recent work analyzes dynamics when agents learn over time (Mandal et al., 2016; Agarwal et al., 2020; Feng et al., 2022). Kong and Schoenebeck (2019) situates multi-task peer prediction in terms of data-processing–monotone information measures, unifying classic mechanisms (Peer Prediction, BTS, CA) and explaining why mechanisms that reward informative agreement discourage uninformative equilibria. On the theoretical front, Schoenebeck and Yu (2020) show that multi-task peer-prediction rules can be learned from data and achieve strong truthfulness, while Zheng et al. (2021) show core limits on what multi-task peer prediction can elicit. Complementary work by Liu and Chen (2017) shows how machine learning can recover the structure needed for peer prediction (“machine-learning aided” elicitation), while Liu and Helmbold (2020) analyze online learning with only peer feedback.
@@
Concurrently, work on LLM-based peer-informed scoring has split into two directions. One line learns a textual scoring rule aligned to a chosen reference label (e.g., an instructor’s grade), fitting to that external signal (Lu et al., 2025); relatedly, Wu and Hartline (2024) scores text against ground-truth instructor reviews via proper scoring rules implemented with LLM oracles. The second line uses an LLM’s token-level likelihoods to compare reports without gold labels—either by predicting a peer’s text or by estimating dependence with peer references (Lu et al., 2024; Xu et al., 2024). By contrast, we target open-web search, where reference labels are unavailable and likelihood-based comparisons across heterogeneous, noisy, and adversarial pages are brittle: we form leave-one-out atomic claims, extract claim-level stances, and score sources by informative peer agreement before re-summarizing.
@@ Appendix B Proofs and Details for Section 2
As justified in the main text, by construction, the held-out set $T_{i}$ is a function of $(q,\boldsymbol{\tau}_{-i})$ only; the scored source $i$ neither selects nor can tailor its content to the realized set. It is therefore natural—and standard in multi-task peer-prediction—to summarize $i$’s behavior on $T_{i}$ by a single set of conditional reporting parameters that do not depend on the claim index $k$. Concretely, conditional on $T_{i}$ there exist constants
@@ Appendix C Equivalence of documents and policies
Appendix D Discussion of Modeling Assumptions
@@
The assumptions mirror standard modeling in the multi-task peer-prediction literature (Shnayder et al., 2016; Dasgupta and Ghosh, 2013; Agarwal et al., 2020). The Leave-One-Out (LOO) construction makes the held-out claim set $T_{i}$ exogenous to the scored source $i$. From source $i$’s perspective, the claims are therefore effectively exchangeable, justifying the use of claim-invariant parameters for $i$ (e.g., $\alpha_{i},\eta_{i}$) while allowing per-claim heterogeneity for its peers. The conditional independence assumption (A2) is the standard separability condition required to identify agreement that is truly informative about the latent ground truth ($\theta_{k}$), as opposed to agreement caused by sources simply copying one another. [^7] While A2 can be violated by near-duplicate sources, this is a known issue that can be effectively mitigated through pre-processing steps like semantic deduplication. Our analysis therefore assumes A2 holds for the set of informationally distinct documents that would remain after such filtering. Finally, the positive peer margin assumption (A3) is weak; it only requires that the peer pool contains some useful signal on average, allowing for some peers or claims to be uninformative.

### v2

@@ 4.2 Results 1: Robustness against adversarial and untruthful sources
As a result of this clear separation between reliable and unereliable sources, we are able to see significant improvement gain in accuracy for both the NQ and ClashEval dataset in Table 1 and 2. Fluency also improves: see App. H.1 (Table 3).
@@
Initial Synthesis 40.8% 26.2% 31.9% 25.1%
Majority Prompt 43.4% 26.5% 32.9% 27.5%
Majority Claims 50.1% 24.7% 33.1% 38.6%
Our Method (TTS) 76.1% 37.2% 50.0% 72.3%
@@
Initial Synthesis 49.3% 22.0% 30.4% 15.6%
Majority Prompt 58.7% 25.4% 35.4% 30.2%
Majority Claims 63.6% 20.4% 30.9% 38.4%
Our Method (TTS) 86.2% 31.5% 56.1% 77.1%
@@
Figure 3: Score separation and incentives. Left: Informative-agreement scores separate reliable from unreliable sources. Right: Truthful behavior is payoff-maximizing against deviations in stance.
@@
Computational Efficiency
With the A/B grouping scheme in Sec. 3.5, the dominant cost of TTS comes from stance extraction over source–claim pairs, which scales linearly in the number of sources and claims and is parallelizable. On our main NQ and ClashEval runs, the average query complexity with 7 retrieved sources is 174,346 input tokens, 12,733 output tokens, and 116,858 thinking tokens across all calls to gemini-2.5-flash-lite.
At current commercial pricing, this costs $\approx$\$0.07 per query, though in-house models would be substantially cheaper. Moreover, TTS need not run on every query; as detailed in Appendix D, it can be applied to sampled traffic to accumulate source reputation signals.
@@ 4.3 Result 2: Robustness against coordinated, uninformative behavior
One of the main advantages of the adapted multi-task peer prediction scoring rule is its robustness to coordinated, uninformative behavior, a canonical failure mode for simpler consensus-based systems. We test this in the ClashEval dataset by introducing a bloc of four “uninformative” sources strategically authored to contradict every claim. As shown in Figure 4, the naive majority-based scoring fails catastrophically. It not only rewards the colluding dummy sources, but as a byproduct, this pollution of the peer pool also falsely elevates the score of the adversarial source, causing it to be ranked higher than the genuinely truthful documents. In contrast, TTS correctly assigns near-zero scores to the uninformative bloc and robustly preserves the correct reliability ranking. More details are given in Appendix H.4.
@@
These empirical findings are consistent with our theoretical guarantees, demonstrating that the TTS framework makes truthful, careful reporting the most effective strategy for a source to be included in the final summary.
@@ Appendix A Related Works
Commercial search has already shifted toward LLM-written overviews that synthesize multiple pages (Google’s AI Overviews; Microsoft’s Copilot Search in Bing, Perplexity AI). In these experiences, citations are shown but the LLM determines salience and framing, moving competition from ranked links to representation in the overview itself [Google Search Blog, , Bing Team, , Perplexity AI, ].
@@
Our setting, multiple web sources of uneven quality, possibly in conflict, aligns with many research papers in the RAG domain. First, recent benchmarks systematize how to stress-test RAG beyond vanilla QA: they evaluate robustness to noise, counterfactuals, and long-context alternatives; provide explainable testbeds and failure analytics; and introduce standard tooling to compare systems [Chen et al., 2024, Friel et al., 2024, Rau et al., 2024, Li et al., 2025]. Building on such evaluations, a second line studies how models arbitrate conflicts between internal priors and external evidence: ClashEval shows that state-of-the-art LLMs frequently adopt incorrect retrieved content over correct priors under controlled perturbations [Wu et al., 2024]; subsequent methods reason explicitly over disagreement, e.g., AstuteRAG which elicits parametric knowledge, clusters internal/external evidence into consistent vs. conflicting sets, and finalizes answers by reliability [Wang et al., 2024a], and MADAM-RAG which assigns each document to an agent, debates, and aggregates, evaluated on the RAMDocs dataset with ambiguity, misinformation, and noise [Wang et al., 2025a]. However, not all search scenarios should (or can) rely on internal priors of LLMs: for breaking news and evolving events, priors are stale. Consistent with our Introduction, we focus on settings where we either omit priors or treat them as just one more source when helpful. A third line introduces self-monitoring and corrective control: Self-RAG learns when to retrieve and to critique its own generations via reflection tokens [Asai et al., 2024], while Corrective RAG (CRAG) adds a retrieval-quality evaluator that triggers fallback actions (broaden web search, decompose/recompose) when evidence appears unreliable [Yan et al., 2024]. Finally, when retrieval itself is corrupted, recent work documents how “blocker” or misleading documents can drive RAG below non-RAG baselines [Zeng et al., 2025] and develops defenses that isolate per-passage influence and certifiably bound the impact of a limited number of corrupted contexts [Xiang et al., 2024]. Complementary audits quantify how small amounts of synthetic misinformation materially degrade knowledge-intensive QA [Pan et al., 2023]. These techniques harden fixed pipelines; by contrast, our goal is to reshape incentives so truthful reporting is the best strategy for sources.
@@
Because summaries now mediate attention, sources adapt to whatever the system rewards. Beyond classical prompt-injection via web content [Greshake et al., 2023], Nestaas et al. [2024] study so-called adversarial search engine optimization (SEO)—deliberately crafting pages to make an LLM favor them regardless of factual merit—including preference-manipulation attacks demonstrated against production LLM search and plugin ecosystems. Reports of hidden instructions in scholarly submissions targeting LLM-assisted review illustrate similar gaming incentives [Gibney, 2025]. Our approach aims to dissuade such user-unfriendly manipulation by changing how sources are scored and fed into the summary.
@@
Incentive-aligned elicitation without verifiable truth is the province of peer prediction. Foundations include the Peer-Prediction method [Miller et al., 2005] and Bayesian Truth Serum (BTS) [Prelec, 2004], with robust BTS variants that work in small populations and for non-binary or continuous signals [Witkowski and Parkes, 2012, Radanovic and Faltings, 2013]. Multi-task mechanisms address effort and uninformative agreement: output-agreement–style rules and their refinements establish strong or informed truthfulness given structure on signals [Dasgupta and Ghosh, 2013, Shnayder et al., 2016]. Of particular relevance is Correlated Agreement (CA), which rewards informative (surprising) agreement across tasks rather than raw consensus; extensions handle heterogeneous tasks and heterogeneous user types, and recent work analyzes dynamics when agents learn over time [Mandal et al., 2016, Agarwal et al., 2020, Feng et al., 2022]. Kong and Schoenebeck [2019] situates multi-task peer prediction in terms of data-processing–monotone information measures, unifying classic mechanisms (Peer Prediction, BTS, CA) and explaining why mechanisms that reward informative agreement discourage uninformative equilibria. On the theoretical front, Schoenebeck and Yu [2020] show that multi-task peer-prediction rules can be learned from data and achieve strong truthfulness, while Zheng et al. [2021] show core limits on what multi-task peer prediction can elicit. Complementary work by Liu and Chen [2017] shows how machine learning can recover the structure needed for peer prediction (“machine-learning aided” elicitation), while Liu and Helmbold [2020] analyze online learning with only peer feedback.
@@
Concurrently, work on LLM-based peer-informed scoring has split into two directions. One line learns a textual scoring rule aligned to a chosen reference label (e.g., an instructor’s grade), fitting to that external signal [Lu et al., 2025]; relatedly, Wu and Hartline [2024] scores text against ground-truth instructor reviews via proper scoring rules implemented with LLM oracles. The second line uses an LLM’s token-level likelihoods to compare reports without gold labels—either by predicting a peer’s text or by estimating dependence with peer references [Lu et al., 2024, Xu et al., 2024]. By contrast, we target open-web search, where reference labels are unavailable and likelihood-based comparisons across heterogeneous, noisy, and adversarial pages are brittle: we form leave-one-out atomic claims, extract claim-level stances, and score sources by informative peer agreement before re-summarizing.
Computational Mechanism Design.
Our work also connects to the broader field of computational mechanism design, which leverages algorithms and optimization to design incentive-compatible rules. A prominent line of research utilizes deep learning to automatically learn optimal auction mechanisms [Curry et al., 2025]. Early architectures like RegretNet [Dütting et al., 2024] train neural networks to maximize revenue while minimizing violations of incentive constraints, yielding mechanisms that are approximately incentive-compatible. Subsequent work, such as GemNet [Wang et al., 2024b], improves upon this by enforcing full incentive compatibility (strategyproofness) within the architecture itself, rather than as a soft penalty. These learning-based approaches have been extended to complex settings including data market design [Ravindranath et al., 2023], combinatorial auctions [Wang et al., 2025b], and auctions with budgets [Feng et al., 2018]. Concurrently, recent research has begun to integrate Large Language Models into mechanism design pipelines [Duetting et al., 2024, Dubey et al., 2024, Huang et al., 2025, Soumalias et al., 2025] and how LLM agents behave in traditional mechanisms [Horton, 2023, Shah et al., 2025]. Beyond auctions, there is growing interest in the intersection of Large Language Models and game theory for modeling strategic environments. For instance, Daskalakis et al. [2024] demonstrate that LLMs can be used to extract formal extensive-form game trees and character incentives directly from narrative texts. Our framework occupies a complementary angle: rather than learning mechanism parameters via gradient descent or extracting existing games from text, we apply a specific game-theoretic structure (multi-task peer prediction) onto the RAG process, creating a new game that incentivizes LLM-generated sources to be truthful.
@@ Appendix B Proofs and Details for Section 2
As justified in the main text, by construction, the held-out set $T_{i}$ is a function of $(q,\boldsymbol{\tau}_{-i})$ only; the scored source $i$ neither selects nor can tailor its content to the realized set. It is therefore natural and standard in multi-task peer-prediction to summarize $i$’s behavior on $T_{i}$ by a single set of conditional reporting parameters that do not depend on the claim index $k$. Concretely, conditional on $T_{i}$ there exist constants
@@ Appendix C Equivalence of documents and policies
Appendix D Discussion of Modeling Assumptions and Future Directions
@@
The assumptions mirror standard modeling in the multi-task peer-prediction literature [Shnayder et al., 2016, Dasgupta and Ghosh, 2013, Agarwal et al., 2020]. The Leave-One-Out (LOO) construction makes the held-out claim set $T_{i}$ exogenous to the scored source $i$. From source $i$’s perspective, the claims are therefore effectively exchangeable, justifying the use of claim-invariant parameters for $i$ (e.g., $\alpha_{i},\eta_{i}$) while allowing per-claim heterogeneity for its peers. The conditional independence assumption (A2) is the standard separability condition required to identify agreement that is truly informative about the latent ground truth ($\theta_{k}$), as opposed to agreement caused by sources simply copying one another. [^7] While A2 can be violated by near-duplicate sources, this is a known issue that can be effectively mitigated through pre-processing steps like semantic deduplication. Our analysis therefore assumes A2 holds for the set of informationally distinct documents that would remain after such filtering. Finally, the positive peer margin assumption (A3) is weak; it only requires that the peer pool contains some useful signal on average, allowing for some peers or claims to be uninformative.
Role in practice and practical desiderata.
These assumptions should be viewed as an idealized lens rather than a literal description of the web. We adopt them for the same reason as prior multi-task peer-prediction work: they give a clean, tractable model in which we can formally relate our peer-consistency score to latent source accuracy and prove that truthful reporting is a best response. In practical systems, the bar is different: we do not need to eliminate all strategic behavior, but to make systematic manipulation significantly more costly and less rewarding. Our experiments already operate in a regime where exchangeability and conditional independence hold only approximately (LLM-generated sources, realistic QA queries, heterogeneous documents), yet TTS still substantially reduces the influence of deceptive or adversarial sources. This suggests that the mechanism can serve as a useful building block for real-world deployments even when the theoretical assumptions are only approximately satisfied.
@@
Discussion on diversity and unique claims.
A natural concern is whether our source selection rule might discourage sources from contributing unique information, thereby harming diversity in the retrieved context and downstream QA.
First, our mechanism does not penalize uniqueness per se. The stance extractor distinguishes between contradiction and mere lack of topic overlap: if a source contributes a claim that is not mentioned elsewhere but does not conflict with the shared core of other sources, it is labeled as abstain rather than contradict. Such unique-but-compatible claims are still eligible for inclusion. In practice, if a source’s overall score $\widehat{w}_{i}$ passes the threshold, its unique claims can be synthesized into the final answer and enrich the summary. A source is not punished simply for saying something additional that others do not say.
The more delicate case is when a unique claim directly contradicts a confident majority—for example, ten sources assert $A$ and one source asserts $B$, with similar apparent quality and $A$ and $B$ mutually exclusive. In our theoretical framework we explicitly assume that we do not have access to ground truth and we do not rely on static priors (e.g., asking the same LLM to arbitrate using its internal world knowledge), which we believe mirrors the regime of rapidly evolving search. In this model, we deliberately restrict ourselves to using only the mutual information between sources: the scoring rule, motivated by multi-task peer prediction, rewards sources that are mutually informative and uses the off-task term to discount uninformative “herd” equilibria. But if all the observable evidence is “10 say $A$, 1 says BB” with no additional prior signal, there is no purely interaction-based reason to systematically favor $B$ over $A$.
Our goal in this paper is to analyze that interaction-based mechanism in isolation, not to design a complete end-to-end retrieval stack. In practical deployments, it is natural—and often desirable—to combine our mechanism with simple prior-based components, and our design is explicitly compatible with that. Two examples are:
  • Reputation weighting as discussed above, where practitioners can assign higher prior weight to known reliable outlets (e.g., major news organizations, official documentation) than to low-credibility sources. This provides a principled and explicit way to inject domain knowledge, so that a single high-reputation dissenting source can outweigh many low-reputation ones if desired.
  • Composability with other routines. Our method is modular and can be applied on top of existing retrieval and filtering steps (e.g., domain whitelists, de-duplication, diversity-promoting selection). In that case, TTS is used to score and prune among sources that have already passed those prior-based filters, rather than acting as the sole decision mechanism.
In summary, (i) unique but non-contradictory claims are not penalized by our stance model and can still improve coverage, and (ii) the treatment of rare “majority vs. minority” conflicts that genuinely require exogenous priors is an application-level design choice. Our mechanism is intended to serve as a building block that provides incentive-aligned aggregation based on cross-source interactions, and can be combined with reputation signals or other system components when richer priors are available.
D.1 Future Directions
Our analysis in the main text and in this appendix is intentionally static and single-query: we model one round of retrieval and summarization, assume fixed prior reliabilities, and study how the mechanism shapes incentives within that round. A natural next step is to extend this to a dynamic, repeated setting. In particular, the reputation-weighted variant in Section D suggests a natural repeated-game view: per-query scores can be aggregated into long-run reputations that influence future exposure and retrieval. Formally analyzing the resulting dynamics—how strategic sources adapt over time and whether truthful equilibria emerge or are stable under best-response dynamics—is an interesting direction for future work.
A second direction is to study TTS more explicitly in multi-hop and long-context regimes. Our mechanism itself only requires (i) the ability to decompose answers into atomic claims and (ii) to extract per-source stances on those claims, the theoretical guarantees carry over as the number and structure of claims become richer. What changes is primarily computational: multi-hop reasoning and longer contexts increase the number of claims and stance evaluations. Exploring hierarchical variants (e.g., scoring coarse-grained claims first and expanding only contentious regions) and more aggressive claim subsampling, and evaluating these variants on dedicated multi-hop / long-context benchmarks, are promising extensions.
Finally, on the empirical side, we see our current synthetic-but-model-generated sources as a controlled, forward-looking proxy for a future incentive environment. A complementary line of work is to (i) simulate more realistic strategic behaviors at scale (e.g., by training or prompting models to optimize explicitly for inclusion under particular overview systems), and (ii) curate or collect real-world datasets where content has been optimized for LLM-based summaries or where adversarial behavior is known to occur. Coupled with richer human evaluation and qualitative analyses, such datasets would allow a more fine-grained assessment of how TTS interacts with real-world manipulative tactics, and how it can be combined with retrieval-stage defenses and other robustness methods in deployed systems, and (iii) combine TTS with existing robustness methodology. TTS targets rapidly developing web search setting where reliable priors are non-existent and it is not possible to utilize an LLM-judge, internal debate, or methodologies from other papers surveyed in App. A. That said, TTS is designed to be a modular source-filtering and reweighting layer: it can be inserted before or alongside retrieval-robust generation procedures to produce an incentive-aligned pool of documents, after which one could still apply Self-RAG, AstuteRAG, debate-style generation, etc. Future work and practical application can seek to combine those aspects to design systems that are robust in all scenarios.